OpenStudy (anonymous):

Evaluate the integral by interpreting it in terms of areas. integral (-1

OpenStudy (anonymous):

\[I=\int\limits_{-1}^{0}\left( 5+\sqrt{1-x^2} \right)dx\] \[=\int\limits_{-1}^{0}5~dx+\int\limits_{-1}^{0}\sqrt{1-x^2}dx\]

OpenStudy (anonymous):

Is the answer -5?

OpenStudy (anonymous):

I found the answer as\[5+\frac{ \pi }{ 4 } \] through correctly graphing. Thanks for the help though @surjithayer

OpenStudy (anonymous):

Yw